OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (anonymous):

First, find the slope, which is \[\frac{ 4-(-4) }{ 1-(-1) }\] which is equal to 4.

OpenStudy (anonymous):

Then, use the point-slope form, y-4=4(x-1), and simplify to get y=4x

OpenStudy (anonymous):

Thank you so much!